Use the following data to find the direct labor efficiency variance if the company produced 3,500 units during the period. Di- Ο $21,000 favorable. Ο $6,125 unfavorable. Ο $6,125 favorable. Ο $7,000 unfavorable. Ο $7,000 favorable.
Answer #1

Direct labor efficiency variance:

= (Actual hours - Standard hours) × Standard rate

Actual hours = 12,250

Standard hours = 3500 × 4 = 14,000

Standard rate = $ 12 per hour

= (12,250 - 14,000) × $ 12

= 1750 × $ 12

Direct labor efficiency ratio = $ 21,000 favorable (AH < SH)